Room temperature is 10 ° C and relative humidity 75%. Temperature raised to 20 ° C. What is the relative humidity of the air now? How much water should be added to the air to get 50% relative humidity? The temperature remains constant.
1
Expert's answer
2019-03-13T12:49:12-0400

Solution:

Relative humidity is the ratio of vapor density to saturation vapor density.

Relative humidity of the air at "10^\circ C:"

"\phi_1=\frac{\rho} {\rho_{01}}100\%\:(1)"


From previous equation we can find vapor density:


"\rho=\frac{\phi_1 \rho_{01}} {100\%}\:(2)"

Saturation vapor density at "10^\circ C\:\:\rho_{01}=9.4\:g/m^3."

Relative humidity of the air at "20^\circ C:"

"\phi_2=\frac{\rho} {\rho_{02}}100\%\:(3)"

Saturation vapor density at "20^\circ C\:\:\rho_{02}=17.3\:g/m^3."

Let's substitute (2) into (3):


"\phi_2=\frac{\phi_1 \rho_{01}100\%} {\rho_{02}100\%}=\frac{\phi_1 \rho_{01}} {\rho_{02}}=\frac{75\%\cdot 9.4} {17.3}=41\%"

Let's find how much water should be added to the air to get 50% relative humidity. Relative humidity of the air in this case can be found using equation:


"\phi_3=\frac{\rho'} {\rho_{03}}100\%\:(4)"

The temperature remains constant, thus "\rho_{03}=\rho_{02}=17.3\:g/m^3."

Using equation (3), we can find vapor density before adding the water:


"\rho=\frac{\phi_2 \rho_{02}} {100\%}\:(5)"


Mass of vapor before adding the water:


"m_1=\rho V=\frac{\phi_2\rho_{02} V} {100\%}\:(6)"

Using equation (4), we can find vapor density after adding the water:


"\rho'=\frac{\phi_3 \rho_{03}} {100\%}\:(7)"

Mass of vapor after adding the water:

"m_2=\rho' V=\frac{\phi_3\rho_{03} V} {100\%}=\frac{\phi_3\rho_{02} V} {100\%}\:(8)"

Mass of water that must be added to the room air:

"\Delta{m}=m_2-m_1=\frac{\rho_{02} V} {100\%} (\phi_3-\phi_2)\:(9)"

Finally:

"\Delta{m}=\frac{17.3 V} {100\%} (50\%-41\%)=1.6V\:(g)"

Answer: "\phi_2=41\%,\:\Delta{m}=1.6V\:g."